# Training-and-prompt-free General Painterly Harmonization Using Image-wise Attention Sharing

### v1

@@
Abstract.
Painterly Image Harmonization aims at seamlessly blending disparate visual elements within a single coherent image. However, previous approaches often encounter significant limitations due to training data constraints, the need for time-consuming fine-tuning, or reliance on additional prompts. To surmount these hurdles, we design a Training-and-prompt-Free General Painterly Harmonization method using image-wise attention sharing (TF-GPH), which integrates a novel “share-attention module”. This module redefines the traditional self-attention mechanism by allowing for comprehensive image-wise attention, facilitating the use of a state-of-the-art pretrained latent diffusion model without the typical training data limitations. Additionally, we further introduce “similarity reweighting” mechanism enhances performance by effectively harnessing cross-image information, surpassing the capabilities of fine-tuning or prompt-based approaches. At last, we recognize the deficiencies in existing benchmarks and propose the ’General Painterly Harmonization Benchmark’, which employs range-based evaluation metrics to more accurately reflect real-world application. Extensive experiments demonstrate the superior efficacy of our method across various benchmarks. The code and web demo are available at https://github.com/BlueDyee/TF-GPH.
diffusion model, attention, image editing, image harmonization, painterly harmonization, style transfer
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32th ACM International Conference on Multimedia; 28 October - 1 November 2024; Melbourne, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Computer visionccs: Computing methodologies Image processingccs: Applied computing Fine arts
@@ Abstract.
Figure 1. An illustrative example showcasing three fundamental tasks in general painterly harmonization: Object Insertion (columns 1 and 2), Object Swapping (columns 3 and 4), and Style Transfer (final column). The top row displays user-created composite images, with the last column’s green box highlighting the background reference. The bottom row reveals the results of our method, adeptly blending the morphology of the inserted object, the semantics of the background and the overall style.
[…]
Teaser.
1. Introduction
Image composition, the art of blending a foreground element from another image with a different background, frequently results in composite images marked by discordant color and illumination between the foreground and background. This discordance can render the composite image inharmonious and unrealistic. To mitigate these issues, image harmonization techniques have been developed, aiming to adjust the foreground appearance to seamlessly blend with its background (Tsai et al., 2017; Lalonde and Efros, 2007; Wu et al., 2019; Tan et al., 2023; Chen et al., 2023; Xing et al., 2022). A specific subset of this field, painterly image harmonization, focuses on the integration of photographic elements into paintings, thereby facilitating artistic edits (Lu et al., 2023a; Cao et al., 2023a; Zhang et al., 2020; Luan et al., 2018; Peng et al., 2019). For example, (Cao et al., 2023a) introduces a novel approach utilizing a dual-domain generator and discriminator to enhance visual harmony across both spatial and frequency domains. Furthermore, (Lu et al., 2023a) unveils PHDiffusion, a model that incorporates a lightweight adaptive encoder and a dual encoder fusion module, striking a better balance between style adaptation and content preservation in the realm of painterly image harmonization.
Despite notable advancements, current painterly image harmonization techniques struggle with generalizability, particularly when confronted with novel art styles or unique content compositions. A promising direction to overcome this limitation is the adoption of one-shot style transfer techniques, as suggested by VCT (Cheng et al., 2023), which facilitate model adaptation through fine-tuning with a single style image. This approach enables networks to reimagine content images by applying stylistic elements learned from the style reference. However, this fine-tuning requirement significantly decreases its practicality and flexibility, as each style requires an additional training process that incurs computational costs 10 to 1000 times higher than a single inference process alone. Currently, the rise of multi-modality models has inspired alternative strategies, such as leveraging text-guided image synthesis to convey style nuances (Lu et al., 2023b; Kwon and Ye, 2022; Hertz et al., 2023). For example, TF-ICON (Lu et al., 2023b) employs carefully crafted text prompts to guide the stylistic integration of objects into their new backgrounds. Yet, the effectiveness of this approach is hindered by the inherent challenge of encapsulating complex visual styles—spanning texture, color gradients, and spatial arrangements—through textual descriptions alone, highlighting the difficulty of translating the multifaceted nature of visual styles into precise and unambiguous textual prompts.
In this work, we propose TF-GPH, a novel diffusion pipeline that operates without the need for additional training or prompts, harnessing the capabilities of the pretrained latent diffusion model (Rombach et al., 2022). Notably, TF-GPH stands as a comprehensive solution, adept at executing a broad spectrum of painterly harmonization tasks, ranging from object insertion and swapping to full style transfers, as demonstrated in Fig. 1. This unified approach not only maintains the computational efficiency of its foundational diffusion pipeline, but also sets new benchmarks, surpassing previous methodologies in diverse evaluations. Specifically, at the core of our innovation is the enhancement of the conventional self-attention mechanism within the diffusion architecture. We introduce a novel share attention layer, engineered to facilitate image-wise rather than patch-limited attention modulation, enabling the composition image to directly attend to both content and style references, eliminating the need of the fine-tuning process to encode style information into diffusion model. Moreover, instead of leveraging the hard-attention adjustments found in prior work (Hertz et al., 2023; Lu et al., 2023b; Gu et al., 2024; Tumanyan et al., 2023; Chefer et al., 2023) to adjust the attention weight toward both content and style references, we further propose the similarity reweighting technique for balancing the attention weight toward both content and style references by scaling the similarity based on given hyperparameters. As such, a controllable harmonic integration of reference image styles can be achieved without the need of additional cross-attention output from prompt. This mechanism gives users the flexibility to subtly influence the intensity of the style, tailoring the output to various artistic preferences by adjusting the hyperparameters.
Finally, a critical challenge in the evaluation of painterly harmonization lies in the limited diversity of test data styles, which are often constrained to the styles seen during training (e.g., the pairing of WikiArt (Tan et al., 2019) with COCO (Lin et al., 2014)). This limitation poorly reflects the broad spectrum of styles encountered in real-world scenarios, such as manga or sumi-e, which typically fall outside the distribution of common datasets. To bridge this gap and more accurately assess the zero-shot capabilities of TF-GPH, we introduce the “General Painterly Harmonization Benchmark” (GPH Benchmark). This new benchmark is designed not only to encompass a trio of harmonization tasks—object insertion, object swapping, and style transfer, but also to include different content and style references, thereby ensuring a robust and comprehensive evaluation framework. The validity and effectiveness of the GPH Benchmark are further substantiated through detailed qualitative and quantitative analysis presented in Sec. 5. Additionally, existing metrics solely emphasize either content or style similarity without adequately addressing user preference, i.e., different balances between stylization and content preservation for different samples. To link quantitative indices with user experience, we advocate for the adoption of range-based metrics. The range-based metrics compute both the lower bound and upper bound of the stylization and content-preservation strength among given data, viewing the applicable range as an indicator of adaptability across various scenarios; the wider range should be able to cater to more niches. Our contributions can be summarized as follows:
  1. (1)
    We introduce the TF-GPH framework, the first training-and-prompt-free pipeline using a diffusion model designed for general painterly harmonization.
  2. (2)
    Our proposed share-attention layer with similarity reweighting not only shows promising results in painterly harmonization tasks, but can serve as a more general form of the self-attention layer, providing a flexible option for attention-based image editing methods. [^1] Due to the space constraint, please refer to the experiments in Appendix H of the supplementary materials.
  3. (3)
    We propose the GPH Benchmark, consisting of various data for real-world usage, together with a range-based metric to align model performance with user experience.
2. Related Work
2.1. Image Harmonization.
Image Harmonization can be divided into two main categories: Realistic Image Harmonization and Painterly Image Harmonization. Realistic Image Harmonization methods (Zhang et al., 2021; Cong et al., 2020; Lin et al., 2018; Azadi et al., 2020; Hong et al., 2022; Jiang et al., 2021; Zhang et al., 2020; Chen et al., 2023; Niu et al., 2021; Xing et al., 2022), focus on blending objects into new backgrounds with an emphasis on consistent illumination, edge alignment, and shadow integrity. On the other hand, Painterly Image Harmonization (Lu et al., 2023a; Cao et al., 2023a; Meng et al., 2022; Tsai et al., 2017; Wu et al., 2019; Peng et al., 2019; Yan et al., 2022), aims to artistically integrate objects into paintings, centering on stylistic congruence rather than realistic alignment, and strives to merge both artistic styles and textures, as seen in GP-GAN (Wu et al., 2019) which innovates by incorporating Poisson gradients into the training loss to enhance style blending. PHDNet (Cao et al., 2023a) advances this by integrating style loss (Gatys et al., 2017) and the AdaIN module (Huang and Belongie, 2017) for more refined style adaptation. Furthermore, PHDiffusion (Lu et al., 2023a) represents a significant pivot by adapting the GAN framework to a diffusion model with an adaptive encoder. Despite their advancements, these methods predominantly rely on extensive training processes, potentially limiting their adaptability to novel styles. In contrast, our proposed method eliminates the need for training, offering direct application to unseen styles, thereby broadening the scope and utility of painterly image harmonization.
2.2. Style Transfer.
Style transfer aims to modify the style of a content image to reflect a specified reference style, with existing methods generally falling into two categories: optimization-based and feedforward-based. Optimization-based techniques, such as (Kwon and Ye, 2022; Gatys et al., 2017; Li et al., 2017), refine the image by aligning it with features extracted from the style reference; for instance, CLIPStyler (Kwon and Ye, 2022) leverages a pre-trained CLIP model (Radford et al., 2021) for this purpose. Conversely, feedforward-based approaches, like StyTr2 (Deng et al., 2022), utilize a Vision Transformer (ViT)-based architecture to reconstruct the content image using style feature patches. Among these, QuantArt (Huang et al., 2023) uniquely employ the vector quantization strategy to finely balance content preservation with style. Additionally, with the evolution of diffusion models, methods like VCT (Cheng et al., 2023) and InST (Zhang et al., 2023) have emerged, advocating for the fine-tuning of generative models with a style reference to embed style directly into the model’s architecture. Our proposed TF-GPH method, in contrast, delivers superior performance across a variety of benchmarks without the necessity for fine-tuning, significantly reducing the time required for style transformation.
2.3. Attention-based Image Editing
Manipulation of self-attention and cross-attention layers within diffusion UNet architectures is a common strategy in contemporary image editing methodologies(Epstein et al., 2023; Tumanyan et al., 2023; Chefer et al., 2023; Gu et al., 2024; Lu et al., 2023b; Hertz et al., 2023). Notable examples include P2P (Hertz et al., 2023), which employs prompt-driven cross-attention to modify objects within images, and PhotoSwap (Gu et al., 2024), which utilizes a fine-tuning approach (Ruiz et al., 2023) to position personal input references effectively . Similarly, TF-ICON (Lu et al., 2023b) integrates both self-attention and cross-attention outputs into its diffusion model to seamlessly integrate objects into various backgrounds without the need for additional training, closely aligning our approach. However, all the methods like P2P, PhotoSwap, and TF-ICON depend on prompts that describe the image, which can pose challenges when suitable prompts are unavailable. In contrast, our TF-GPH method eliminates the need for such prompts, relying solely on image inputs and thus enhancing its applicability to a broader range of scenarios.
@@ 2.3. Attention-based Image Editing
Figure 2. The architecture of our proposed TF-GPH method involves several stages. Initially, we feed the denoising U-Net with the inverse latent $Z_{t}$, and during the first $l<L_{\text{share}}-1$ layers of the U-Net, the three latent representations, $z^{\text{f}}_{t}$, $z^{\text{b}}_{t}$, and $z^{\text{comp}}_{t}$, are forwarded separately to the Attention Module. Afterward, they are fed into the Share-Attention Module (the blue part below), obtaining their image-wise attention via Eq. (5). In the end, the output harmonized image $I^{\text{o}}$ is produced.
3. Preliminary
3.1. Diffusion Probabilistic Models
Our approach leverages advances in the forefront of image generation technology, specifically through the implementation of diffusion probabilistic models (DPMs) (Rombach et al., 2022; Dhariwal and Nichol, 2021; Ho et al., 2020; Nichol and Dhariwal, 2021; Song et al., 2021a). The underlying concept behind DPMs is straightforward but powerful: it systematically reduces noise within an image in a stepwise manner. Within the framework of score-based methodologies(Kingma et al., 2021; Song et al., 2021b; Lu et al., 2022), the diffusion process for a specific time step $t\in[0,T]$ can be encapsulated by the following stochastic differential equation (SDE):
(1) $\text{d}x_{t}=f(x_{t},t)\text{d}t+g(t)\text{d}w_{t},$
where $w_{t}$ denotes the standard Wiener process, $f(x_{t},t)$ is identified as the drift coefficient describing the distribution shift and $g(t)$ represents the diffusion coefficient that modulates the intensity of the noise. To counteract the initial noise and restore the image to its original state, a reverse-time SDE is employed:
(2) $\text{d}x_{t}=\left[f(x_{t},t)-g(t)^{2}\nabla\text{log}p_{t}(x_{t})\right]\text{d}t+g(t)\text{d}\bar{w}_{t},$
where $\nabla\text{log}p_{t}(x_{t})$ is the scoring function, often approximated by a neural network, denoted as $\epsilon_{\theta}(x_{t},t)$. The model parameters, represented by $\theta$, are optimized through a noise reduction training objective (Song et al., 2021b; Ho et al., 2020), which facilitates the conversion of noisy inputs back to their original noise-free counterparts. Throughout this paper, we follow the Variance Preserving SDE (Song et al., 2021b) form.
3.2. DPM Inversion
In the domain of diffusion-based image editing (Gu et al., 2024; Liew et al., 2022; Cao et al., 2023b; Shi et al., 2023; Couairon et al., 2023; Kwon et al., 2022; Lu et al., 2023b; Hertz et al., 2023), inversion plays a crucial role. Successful inversion signifies the model’s capability to regenerate a reference image, thereby enabling modifications in the regenerated image to effectuate property alterations in the edited output. Conventional approaches predominantly utilize the DDIM (Song et al., 2021a) to reversibly return the image to its noise form through an ordinary differential equation:
(3) $\text{d}x_{t}=\left[f(x_{t},t)-g(t)^{2}\nabla\text{log}p_{t}(x_{t})\right]\text{d}t.$
However, the efficacy of DDIM diminishes when the images is out of distribution, necessitating additional fine-tuning (Ruiz et al., 2023; Gu et al., 2024; Shi et al., 2023). Addressing this limitation, DPM-Solver++ (Lu et al., 2022) emerges as a solution, which we adopt for reconstruction, adept at inverting images across a wider array of distributions without the prerequisite of fine-tuning(Ruiz et al., 2023; Gu et al., 2024; Shi et al., 2023), even in scenarios involving personalized input. In order to improve the stability of the inversion process, TF-ICON (Lu et al., 2023b) introduces the use of an exceptional prompt embedding $\rho_{\text{exceptional}}$, which differs from the conventional null prompt $\phi$, aligning more congruently with the mechanics of native generation.
4. Method
@@
Our research aims to facilitate a general form of painterly harmonization based on images only without the additional need for prompts, which can facilitate a variety of applications, i.e., the insertion of objects, object swapping, and style transfer. Formally, given a background painting $I^{\text{b}}$, a foreground object image $I^{\text{f}}$, and $I^{\text{comp}}$, which is the user-specified composition that guides the size and position of the foreground object on the background painting, the goal of painterly harmonization is to transfer the style from $I^{\text{b}}$ to the object in $I^{\text{comp}}$ seamlessly, resulting a harmonized image $I^{\text{o}}$.
@@
To address the challenge of painterly harmonization, we introduce a novel framework titled Training-and-Prompt-Free General Painterly Harmonization using a Diffusion Model (TF-GPH), as depicted in Fig. 2. Our method ingeniously leverages a diffusion model to harmonize elements within a painterly composition without the need for explicit training or user-defined prompts. Specifically, the inputs—foreground $I^{\text{f}}$, background $I^{\text{b}}$, and composite $I^{\text{comp}}$-are initially processed through an inversion mechanism equipped with either a standard null prompt embedding or a predefined exceptional prompt embedding $\rho_{\text{exceptional}}$, which has demonstrated its ability for stabilizing inversion process (Lu et al., 2023b). Subsequently, a denoising operation is applied concurrently to all three images, during which the composite image $I^{\text{comp}}$ is enriched with style attributes.
The core of our architecture is the Shared Attention Module, a novel component designed to extract and amalgamate the foreground object details from $I^{\text{f}}$ with the stylistic elements of $I^{\text{b}}$, ensuring the coherence of the final output $I^{\text{o}}$. This integration is further refined through our Similarity Reweighting technique, which optimizes the attention matrix to resonate with the underlying feature distribution more effectively. Although our framework is versatile enough to support both object swapping and style transfer—with the former viewed as a semantically richer variant of object insertion and the latter as a broader aspect of the same—our discussion here is confined to object insertion for brevity.
4.1. Share-Attention Module
@@
In the framework of diffusion models, the attention mechanisms (Vaswani et al., 2017) are essential to capture characteristic details, facilitating both the elimination of noise and the enhancement of context information. Specifically, the self-attention module plays a vital role in synthesizing the output by internalizing and using the inherent data characteristics, while the cross-attention module is instrumental in incorporating contextual information from various modalities, e.g., images, text, and audio, thus amplifying the conditional impact on the resultant images. Since our approach does not need an additional prompt to guide the fusion, we can simply utilize self-attention to ensure that the background style is harmonically fused into the composition image during the denoising process.
@@
Given three input images $I^{\text{f}}$, $I^{\text{b}}$, and $I^{\text{comp}}$, these images are first compressed by a VAE encoder (Rombach et al., 2022) into latent representations $z^{\text{f}}_{0}\in\mathbb{R}^{w\times h\times d}$, $z^{\text{b}}_{0}\in\mathbb{R}^{w\times h\times d}$, $z^{\text{comp}}_{0}\in\mathbb{R}^{w\times h\times d}$, respectively, where $w$ and $h$ denote the width and height of the latent shape, $d$ is the feature channels and the subscript 00 denotes the initial timestep of the diffusion process. Next, we apply the DPM-Solver++ inversion to convert the initial latents $z^{\text{f}}_{0}$, $z^{\text{b}}_{0}$, and $z^{\text{comp}}_{0}$ to noisy latents $z^{\text{f}}_{T}$, $z^{\text{b}}_{T}$, and $z^{\text{comp}}_{T}$, where $z_{T}\sim N(0;\tilde{\sigma}_{T}^{2})$. This preprocess enabling the image modification during subsequent reconstruction process.
@@
During the reconstruction process from the time step $T$ to 00, we incorporate the style feature into $z^{\text{comp}}_{t}$ using our proposed shared attention module. This module can be viewed as a more general form of the self-attention module, allowing for feature flow between images within the same batch. Specifically, the traditional self-attention module projects the input feature $z\in\mathbb{R}^{m\times d}$ of length $m=(w\cdot h)$ onto the corresponding $Q,K,V\in\mathbb{R}^{m\times d}$ through learned linear layers inside the original self-attention module and computes the attention matrix $A\in\mathbb{R}^{m\times d}$ as follows.
(4) $A(Q,K,V)=\text{Softmax}\left(QK^{T}/\sqrt{d}\right)V,$
@@
To enable the flow of feature information between images during the attention operation, we should consider three images at the same time instead of processing each attention matrix independently. To create the query, key, and value from three different inputs, we first concatenate three input latents on the first dimension to form $Z_{T}\in\mathbb{R}^{(3\cdot m)\times d}$ = [$z^{\text{f}}_{T}$, $z^{\text{b}}_{T}$, $z^{\text{comp}}_{T}$]. Then we project the latent $Z_{T}$ into the corresponding $\hat{Q},\hat{K},\hat{V}\in\mathbb{R}^{(3\cdot m)\times d}$. However, directly feeding $\hat{Q},\hat{K},\hat{V}$ into Eq. (4) may disrupt the content of $z^{\text{b}}$ and $z^{\text{b}}$ since the additional attention from other images making the latent differ from original reconstruction without attention from others. To keep the content of $z^{\text{f}}$ and $z^{\text{b}}$ intact for correctly guiding the harmonization of $z^{\text{comp}}$, we propose a specially designed mask $\hat{M}\in\mathbb{R}^{(3\cdot m)\times(3\cdot m)}$ that allows $z^{\text{comp}}$ to utilize information from $z^{\text{f}}$ and $z^{\text{b}}$ while keeping $z^{\text{f}}$ and $z^{\text{b}}$ intact. The shared attention equation is thus calculated by:
(5) $\hat{A}(\hat{Q},\hat{K},\hat{V})=\text{Softmax}\left(\hat{M}\odot(\hat{Q}\hat{K}^{T})/\sqrt{d}\right)\hat{V},$

### v2

# Training-and-Prompt-Free General Painterly Harmonization via Zero-Shot Disentenglement on Style and Content References

@@
Abstract
Painterly image harmonization aims at seamlessly blending disparate visual elements within a single image. However, previous approaches often struggle due to limitations in training data or reliance on additional prompts, leading to inharmonious and content-disrupted output. To surmount these hurdles, we design a Training-and-prompt-Free General Painterly Harmonization method (TF-GPH). TF-GPH incorporates a novel “Similarity Disentangle Mask”, which disentangles the foreground content and background image by redirecting their attention to corresponding reference images, enhancing the attention mechanism for multi-image inputs. Additionally, we propose a “Similarity Reweighting” mechanism to balance harmonization between stylization and content preservation. This mechanism minimizes content disruption by prioritizing the content-similar features within the given background style reference. Finally, we address the deficiencies in existing benchmarks by proposing novel range-based evaluation metrics and a new benchmark to better reflect real-world applications. Extensive experiments demonstrate the efficacy of our method in all benchmarks. More detailed in https://github.com/BlueDyee/TF-GPH.
Introduction
Image composition, which involves blending a foreground element from one image with a different background, often results in composite images with mismatched colors and illumination between the foreground and background. Image harmonization techniques have been developed to adjust the appearance of foreground for a seamless integration with the background (Tsai et al. 2017; Wu et al. 2019; Tan et al. 2023; Xing et al. 2022). A specialized area within this field, painterly image harmonization, focuses on integrating elements into paintings to enable artistic edits (Lu et al. 2023; Luan et al. 2018). For instance, ProPIH (Niu et al. 2024b), pioneers progressive painterly harmonization, training the model with different levels of harmonization, enhancing its applicability to real-world scenarios.
Figure 1: Our method overcomes the resolution and staged-progressive painterly harmonization limitations present in the SOTA method ProPIH (Niu et al. 2024b), where users are restricted to selecting stylization strength from one of four stages. In contrast, our approach offer continuously adjustable hyperparameters, allowing for more flexible stylization. Additionally, our method effectively mitigates content disruption issues, such as facial alterations, commonly seen in image-editing methods like ZSTAR (Deng et al. 2023).
Despite notable advancements, current painterly image harmonization techniques still face challenges with generalizability, particularly when dealing with novel art styles or unique content compositions. One promising solution is to leverage insights from other image-editing methods. For instance, (Zhang et al. 2023; Cheng et al. 2023) suggest fine-tuning models to adapt to input styles. However, each styles require additional computational costs that are 10x times higher than a single inference. Alternatively, (Lu, Liu, and Kong 2023; Kwon and Ye 2022) propose text-guided editing strategies, but these approaches are limited by the difficulty of adequately describing complex visual styles through text alone. Recently, methods such as (Cao et al. 2023; Deng et al. 2023) explore training-free techniques. These methods utilize attention-sharing across images combined with techniques like AdaIN(Huang and Belongie 2017) to align content features with style references. While effective, this brute alignment lead to content disruption as shown in Fig. 1.
@@ Introduction
Figure 2: An example demonstrates three tasks in general painterly harmonization: Object Insertion (columns 1 to 3), Object Swapping (columns 4 and 5), and Style Transfer (columns 6 and 7). The top row features user-generated composite images, where green boxes highlighting the style reference of final two. The bottom row showcases the results using our method.
In this work, we present TF-GPH, an innovative diffusion pipeline that operates without additional training or prompts by leveraging the pretrained diffusion model (Rombach et al. 2022). TF-GPH solves a wider range of painterly harmonization tasks, including object insertion, swapping, and style transfers, as illustrated in Fig. 2. In our pipeline, we modify the self-attention mechanism of the diffusion model and adopt the shared attention layer from (Hertz et al. 2024) to enable multi-image input (foreground content reference, background style reference, and composite image). While the shared attention layer can merge similar features across images, such as consecutive frames, the strong self-similarity of the composite image causes it to overlook dissimilar references in this task. To address this, we introduce a novel similarity disentangle mask within the shared attention layer. This mask applied before the softmax operation decouples the foreground and background features by redirecting the composite image’s self-attention to the two reference images. This approach allows precise control over the foreground and background within the composite image by adjusting the attention to their respective references.
Moreover, to address the content disruptions caused by attention adjustments such as addition and AdaIN, as observed in prior work, we propose a similarity-based editing method termed similarity reweighting. This approach balances attention between content and style references by scaling similarity based on user specified hyperparameters. By prioritizing style features that closely match the content features the content disruption thus minimized. By integrating these two aforementioned adjustments into the existing image generation pipeline, we are able to perform image harmonization without requiring additional prompts or training. Additionally, this mechanism offering flexibility to tailor the output continuously from style-free to style-heavy, thereby accommodating various artistic preferences, as shown in Fig. 1.
Finally, a challenge in evaluating painterly harmonization is the limited diversity of test data styles (Tan et al. 2019), which are often restricted to those seen during training. This limitation fails to capture the wide variety of styles encountered in real-world scenarios, such as manga or cartoon. To address this issue, we introduce the “General Painterly Harmonization Benchmark” (GPH Benchmark). This benchmark encompasses three harmonization tasks—object insertion, object swapping, and style transfer—while incorporating diverse content and style references to ensure a comprehensive evaluation. Furthermore, existing metrics typically focus on either content or style similarity without adequately reflecting user preferences for different balances between stylization and content preservation. To bridge this gap, we propose range-based metrics, evaluating both the lower and upper bounds of stylization and content-preservation strength across the dataset. A wider range indicates greater flexibility and the adaptability to various scenarios.
Our contributions can be summarized as follows. 1) We introduce the TF-GPH framework, the first training-and-prompt-free pipeline using a diffusion model designed for general painterly harmonization. 2) Our proposed similarity disentangle mask with similarity reweighting not only shows promising results in painterly harmonization tasks, but also solves the content disruption issue of existing attention-based editing method. 3)We propose the GPH Benchmark, consisting of various data for real-world usage, together with a range-based metric to align model performance with user experience.
Related Work
Image Harmonization
Image Harmonization can be categorized into two main types: Realistic Harmonization and Painterly Harmonization. The former (Zhang et al. 2021; Cong et al. 2020; Lin et al. 2018; Chen et al. 2023) focus on seamlessly integrating objects into new backgrounds with consistent illumination, edge alignment, and shadow integrity. In contrast, Painterly Harmonization (Lu et al. 2023; Cao, Hong, and Niu 2023; Wu et al. 2019) aims to artistically blend objects into paintings, prioritizing stylistic coherence. Recently, ArtoPIH (Niu et al. 2024a) propose learning from painterly objects by using annotated objects within paintings as training data. Additionally, ProPIH (Niu et al. 2024b) introduce a progressive learning approach, improving practical applicability. Despite these advancements, existing methods require training, which can limit their usability. Our proposed method, however, eliminates the need for training, enabling direct application to unseen styles and significantly enhancing the versatility of painterly image harmonization.
Attention-based Image Editing
Manipulation of attention layers within diffusion UNet architectures is a prevalent strategy in modern image editing techniques (Tumanyan et al. 2023; Chefer et al. 2023; Gu et al. 2024; Lu, Liu, and Kong 2023; Hertz et al. 2023). For instance, P2P (Hertz et al. 2023) utilizes prompt-driven cross-attention to modify images, while TF-ICON (Lu, Liu, and Kong 2023) integrates objects into backgrounds by constraining self-attention and cross-attention outputs with given mask. Despite their effectiveness, the reliance on descriptive prompts can be problematic when suitable prompts are not available. In contrast, our TF-GPH method function solely with image inputs, eliminating the need of prompts.
Style Transfer
Style transfer aims to alter the style of a content image to match a specified style. Existing methods generally categorized into optimization-based and feedforward-based approaches. The former (Gatys et al. 2017; Li et al. 2017), refine the image by aligning it with features extracted from the style reference. For example, (Kwon and Ye 2022) utilizes a pre-trained CLIP model (Radford et al. 2021) for this purpose. In contrast, feedforward-based (Deng et al. 2022; Huang et al. 2023) involving VCT (Cheng et al. 2023) and InST (Zhang et al. 2023), which fine-tune models to integrate style into the model’s architecture. Recently, attention-based techniques have been incorporated. For instance, the shared attention module introduced in (Hertz et al. 2024; Deng et al. 2023; Chung, Hyun, and Heo 2024) produces feature-consistent images by sharing attention across multiple images. However, these methods often suffer from content disruption due to the blending of unrelated features from different references. In contrast, TF-GPH minimizes content disruption and achieves superior performance across styles.
@@
Figure 3: The architecture of our proposed TF-GPH method involves several stages. Initially, we feed the denoising U-Net with the inverse latent $Z_{t}$, and during the first $l<L_{\text{share}}-1$ layers of the U-Net, the three latent representations, $z^{\text{f}}_{t}$, $z^{\text{b}}_{t}$, and $z^{\text{c}}_{t}$, are forwarded separately to the Attention Module. Afterward, they are fed into the Share-Attention Module (the blue part below), obtaining their image-wise attention via Eq. (2). In the end, the output harmonized image $I^{\text{o}}$ is produced.
Method
@@
Our research aims to facilitate a general form of painterly harmonization based on images only without the additional need for prompts, which can facilitate various applications, i.e., object insertion, object swapping, and style transfer. Formally, given a foreground object image $I^{\text{f}}$, a background painting $I^{\text{b}}$, and $I^{\text{c}}$, which is the user-specified composition that guides the size and position of the foreground object on the background painting, the goal of painterly harmonization is to transfer the style from $I^{\text{b}}$ to the object from $I^{\text{f}}$ in $I^{\text{c}}$ seamlessly, resulting a harmonized image $I^{\text{o}}$.
@@
To address the challenge of painterly harmonization, we introduce a novel framework titled Training-and-Prompt-Free General Painterly Harmonization (TF-GPH), as depicted in Fig. 3. Specifically, the inputs—foreground $I^{\text{f}}$, background $I^{\text{b}}$, and composite $I^{\text{c}}$-are initially processed through an inversion mechanism equipped with either a null prompt embedding or a exceptional prompt embedding $\rho_{\text{exceptional}}$, which has demonstrated its ability for stabilizing inversion process (Lu, Liu, and Kong 2023). Subsequently, a denoising operation is applied concurrently to all three images, during which the composite image $I^{\text{c}}$ is enriched with style attributes, producing harmonized output $I^{\text{o}}$
The core of our architecture is the Similarity Disentangle Mask, a novel attention mask designed to disentangle the features of the foreground object from the background image of $I^{\text{c}}$ and link them to their corresponding references $I^{\text{f}}$ and $I^{\text{b}}$. After disentanglement, we enhance the influence of the background style reference $I^{\text{b}}$ on the pasted object through our Similarity Reweighting technique. This approach differs from existing attention-based editing techniques, which directly add (Hertz et al. 2023) or adjust the mean/variance of features (Hertz et al. 2024; Chung, Hyun, and Heo 2024), introducing disruption on semantic and structural details. By adjusting the similarity solely, we can minimize content disruption while applying the stylization effect, producing the final painterly harmonized output image $I^{\text{o}}$. Additionally, our framework is versatile enough to support not only painterly harmonization for object insertion—a traditional task of painterly harmonization—but also object swapping and style transfer. The former is viewed as a semantically richer variant of object insertion, and the latter as a broader aspect of the same. We summarize these related tasks under the term “General Painterly Harmonization”.
Attention-based Diffusion UNet
@@
In the framework of diffusion models, the attention mechanisms (Vaswani et al. 2017) are essential to capture characteristic details, facilitating both the elimination of noise and the enhancement of context information. Specifically, the self-attention module plays a vital role in synthesizing the output by internalizing the inherent data characteristics, while the cross-attention module is instrumental in incorporating contextual information from various modalities, e.g. text and audio, thus amplifying the conditional impact on the resultant images. Since our approach does not need an additional prompt to guide the fusion, we can simply utilize self-attention to ensure that the background style is harmonically fused into the composition image during denoising.
@@
Given three input images $I^{\text{f}}$, $I^{\text{b}}$, and $I^{\text{c}}$, these images are first compressed by a VAE encoder (Rombach et al. 2022) into latent representations $z^{\text{f}}_{0}\in\mathbb{R}^{w\times h\times d}$, $z^{\text{b}}_{0}\in\mathbb{R}^{w\times h\times d}$, $z^{\text{c}}_{0}\in\mathbb{R}^{w\times h\times d}$, respectively, where $w$ and $h$ denote the width and height of the latent shape, $d$ is the feature channels and the subscript 00 denotes the initial timestep of the diffusion process. Next, we apply the DPM-Solver++ inversion to convert the initial latents $z^{\text{f}}_{0}$, $z^{\text{b}}_{0}$, and $z^{\text{c}}_{0}$ to noisy latents $z^{\text{f}}_{T}$, $z^{\text{b}}_{T}$, and $z^{\text{c}}_{T}$. This preprocess enabling the image modification during subsequent reconstruction process.
Share-Attention Module
@@
During the reconstruction process from the time step $T$ to 00, we incorporate the style feature into $z^{\text{c}}_{t}$ using the shared attention module. This module can be viewed as a more general form of the self-attention module, allowing for feature flow between input images. Specifically, the traditional self-attention module projects the input feature $z\in\mathbb{R}^{m\times d}$ of length $m=(w\cdot h)$ onto the corresponding $Q,K,V\in\mathbb{R}^{m\times d}$ through learned linear layers inside the original self-attention module and computes the attention matrix $A\in\mathbb{R}^{m\times d}$ as follows.
$A(Q,K,V)=\text{Softmax}\left(QK^{T}/\sqrt{d}\right)V,$ (1)
@@
To enable the flow of feature information between images during the attention operation, we should consider three images at the same time instead of processing each attention matrix independently. To create the query, key, and value from three different inputs, we first concatenate three input latents on the first dimension to form $Z_{T}\in\mathbb{R}^{(3\cdot m)\times d}$ = [$z^{\text{f}}_{T}$, $z^{\text{b}}_{T}$, $z^{\text{c}}_{T}$]. Then we project the latent $Z_{T}$ into the corresponding $\hat{Q},\hat{K},\hat{V}\in\mathbb{R}^{(3\cdot m)\times d}$.
Similarity Disentangle Mask
However, directly feeding $\hat{Q},\hat{K},\hat{V}$ into Eq. (1) may disrupt the features of $z^{\text{f}}$ and $z^{\text{b}}$ since the additional attention from other images making the latent differ from original reconstruction without attention from others. To keep the content of $z^{\text{f}}$ and $z^{\text{b}}$ intact for correctly guiding the harmonization of $z^{\text{c}}$, we propose a specially designed mask called similarity disentangle mask $\hat{M}\in\mathbb{R}^{(3\cdot m)\times(3\cdot m)}$ that allows $z^{\text{c}}$ to utilize information from $z^{\text{f}}$ and $z^{\text{b}}$ while keeping $z^{\text{f}}$ and $z^{\text{b}}$ intact. The shared attention equation is thus calculated by:
$\hat{A}(\hat{Q},\hat{K},\hat{V})=\text{Softmax}\left(\hat{M}\odot(\hat{Q}\hat{K}^{T})/\sqrt{d}\right)\hat{V},$ (2)